# Tac-Man: Tactile-Informed Prior-Free Manipulation of Articulated Objects

### v1

@@ VII Conclusion
The efficacy of Tac-Man is demonstrated through a series of experiments and simulations, highlighting its robustness in diverse scenarios. This includes its ability to handle ambiguous perceptions, adapt to imperfect parameter estimates, and respond effectively to unforeseen perturbations. The versatility of Tac-Man is further evidenced by its ability to manage objects with intricate trajectories, showcasing its practical utility in real-world applications.
In summary, this research makes a timely contribution to the field of robotic manipulation. It introduces a flexible and intuitive solution for interactions with articulated objects in environments marked by complexity and uncertainty. The adoption of a prior-free methodology marks a step forward in developing robotic systems that are more adaptable and capable of performing sophisticated manipulation tasks, paving the way for enhanced automation and efficiency in various applications.
@@ Acknowledgment
Our sincere thanks go to Miss Yujie Fan and Miss Zhen Chen (BIGAI) for their expertise and creativity in figure design. We appreciate the technical support provided by Hongjie Li (PKU) in setting up the simulations. Special acknowledgment is due to Dr. Chi Zhang (BIGAI) and Muzhi Han (UCLA) for their participation in insightful discussions that enriched this work. We are grateful to Yida Niu (PKU) for developing the initial version of the Kinova driver, an essential component of our research setup. Lastly, we thank Mr. Mish Toszeghi (QMUL) for his meticulous proofreading of our manuscript.
-A Material List
For detailed demonstrations related to this study, a comprehensive list of supplementary video URLs is provided. These videos visualize experiments and simulations discussed in this paper. Please refer to Tab. A1 for the complete set of links to these supplementary materials.
TABLE A1: Complete URLs of supplementary materials.
Material Description and URL
Supp. Video S0 Manipulation trajectories in LABEL:fig:teaser
https://vimeo.com/907222643/9793e630b0
Supp. Video S1 Manipulation under ambiguous priors
https://vimeo.com/907224197/f38b1dea62
Supp. Video S2 Manipulation under imperfect priors
https://vimeo.com/907224440/b059314753
Supp. Video S3 Manipulation under unknown priors
https://vimeo.com/907224671/cb9aa51fcb
Supp. Video S4 Manipulation under obsolescent priors
https://vimeo.com/907224990/3b4181a4e0
Supp. Video S5 Simulation studies
https://vimeo.com/907225198/b641a5a84d
-B Details in Simulation Setup
Tac-Man leverages the compliance of the silica gel material on the gripper to track contact. Simulating such compliance within a rigid body simulator presents unique challenges, for which we have implemented a series of specialized designs. These designs are crucial to accurately compute the contact dynamics and effectively mimic the compliance of the gripper’s material. The simulation setup aims to replicate the tactile feedback mechanism as closely as possible to real-world conditions. Key hyper-parameters play a significant role in this process, and their values are determined based on empirical observations and testing. The specific values used in our simulations, which are instrumental in achieving realistic contact dynamics, are enumerated in Tab. A2. This table provides a detailed overview of the hyper-parameters, ensuring transparency and reproducibility of the simulation results.
TABLE A2: Hyper-parameters defined in Secs. -B and -C.
Hyper-parameter $n_{\mathrm{res}}$ $\epsilon^{\mathrm{(sim)}}$ $\zeta_{n}$ $n_{\mathrm{ctrl}}$ $H$ $W$ $w$ $r_{\mathrm{pin}}$ $\eta$
Value 10 $0.00025~{}\mathrm{m}$ $2~{}\mathrm{mm}$ 3, 4, 5, 6 $40~{}\mathrm{cm}$ $60~{}\mathrm{cm}$ $4~{}\mathrm{cm}$ $1.5~{}\mathrm{cm}$ 0.02
Establishing contact
In the simulated environment, a point cloud grid of resolution $n_{\mathrm{res}}\times n_{\mathrm{res}}$ is sampled on each fingertip of the gripper, resulting in $C^{\mathrm{(sim)}}_{0}=\{\mathbf{p}^{0}_{i}|i=1,\dots,m\}$. The notation ($(\cdot)^{\mathrm{(sim)}}$) indicates simulated quantities. To mimic real-world sensor noise, Gaussian noise $\epsilon\sim\mathcal{N}(0,\zeta^{2})$ is added to the point cloud. Once the gripper grasps the handle, we define the initial contact set $C^{\mathrm{(sim)}}_{1}$ as:
$C^{\mathrm{(sim)}}_{1}=\{\mathbf{p}_{i}|\mathcal{D}(\mathbf{p}^{0}_{i})<\epsilon^{\mathrm{(sim)}}\},$ (A1)
where $\mathcal{D}(\mathbf{p})$ denotes the distance from point $\mathbf{p}$ to the gripper’s surface.
Simulating contact
@@
With $C^{\mathrm{(sim)}}_{1}$ established, this contact set is then affixed to the handle’s frame. As the manipulation progresses, $C^{\mathrm{(sim)}}_{1}$ is transformed back to the gripper’s frame $T_{i}$ to update the contact $C^{\mathrm{(sim)}}_{i}$. The iterative process between the execution and recovery stages in the simulation adheres to the methodology detailed in Sec. III and utilizes the same hyper-parameters as in the real-world experiments (Tab. I).
Simulating gripper compliance
To realistically simulate gripper compliance within a rigid body simulation framework, we introduce a mechanism that fixes the joint on the object during the recovery stage. This approach is designed to approximate the flexibility and adaptive nature of an actual gripper’s material, ensuring that the simulated interaction dynamics closely resemble those observed in real-world scenarios.
-C Preprocess PartNet-Mobility
@@
The PartNet-Mobility dataset [36] comprises 1,045 articulated objects, and GAPartNet [38] enriches it with detailed part-level annotations. In this dataset, each object includes a base link and one or more moving links, connected via prismatic or revolute joints. Objects with multiple joints often necessitate sequential manipulations to fully access all their links (for instance, opening an oven door before pulling out the grill inside). For the sake of simplicity in our study, we focus on objects with a single accessible handle. We select objects that necessitate manipulating just one joint by grasping the linked handle. From this refined subset, we randomly chose 50 objects with revolute joints and another 50 with prismatic joints, adhering to these specified criteria.
-D Generate Random 1-DoF Playboards
@@
Our simulation setup, as depicted in Fig. 11(b), incorporates playboards consisting of a baseboard (colored white) and a toy train (colored red). The train’s movement is restricted to a predetermined trajectory (colored grey). To achieve this, we constructed a groove on a flat cube board and set up a distance joint between the board and the toy train, limiting its movement to within the plane of the board. Two cylindrical locating pins beneath the train’s collision model are inserted into the groove, further guiding the train along the set trajectory.
The trajectories are generated as Bézier curves with $n_{\mathrm{ctrl}}$ control points, $B(s),s\in[0,1]$, resulting in $(n_{\mathrm{ctrl}}-1)$-order curves. The control points are uniformly sampled and scaled to fit within a $H\mathrm{cm}\times W\mathrm{cm}$ area. This scaling ensures that the trajectories, confined within the convex hull of the control points, remain within the defined area. Any self-intersecting curves, taking into account a padding of $w$, are discarded. The groove, based on these curves, is carved into a $(H+2w)\mathrm{cm}\times(W+2w)\mathrm{cm}$ board. The initial position of the toy train is set at $B(\eta)$, where $\eta$ is a small positive value to prevent initial pin-border collisions. Fig. 10(b) illustrates an example curve and corresponding play-board, with control points marked as red ‘x’s. Additional examples in Fig. 10(c) demonstrate that higher-order curves generally exhibit greater complexity, posing increased manipulation challenges. The hyper-parameters for this setup are listed in Tab. A2.

### v3

@@ VII Conclusion
The efficacy of Tac-Man is demonstrated through a series of experiments and simulations, highlighting its robustness in diverse scenarios. This includes its ability to handle ambiguous perceptions, adapt to imperfect parameter estimates, and respond effectively to unforeseen perturbations. The versatility of Tac-Man is further evidenced by its ability to manage objects with intricate trajectories in simulation.
In summary, this research makes a timely contribution to the field of robotic manipulation. It introduces a flexible and intuitive solution for interactions with articulated objects in environments marked by complexity and uncertainty. The adoption of a prior-free methodology marks a step towards robotic systems that are more adaptable and capable of performing sophisticated manipulation tasks, paving the way for enhanced automation and efficiency in various applications.
@@ Acknowledgment
We thank Hongjie Li (PKU), Yida Niu (PKU), and Zimo He (PKU) for setting up the experiments; Prof. Hao Dong (PKU), Zeyi Li (PKU), and Ruihai Wu (PKU) for setting up the Franka Robot; Dr. Chi Zhang (BIGAI) and Dr. Muzhi Han (UCLA) for insightful discussions; and Mr. Mish Toszeghi (QMUL) for meticulously proofreading our manuscript.
-A Details in Simulation Setup
TABLE A1: Hyper-parameters defined in Secs. -A and -B.
Para. $n_{\mathrm{res}}$ $\prescript{\mathrm{sim}}{}{\epsilon}$ $\zeta_{n}$ $n_{\mathrm{ctrl}}$ $H$ $W$ $w$ $r_{\mathrm{pin}}$ $\eta$
Value $10$ $0.25~{}\mathrm{mm}$ $2~{}\mathrm{mm}$ $3$-$6$ $40~{}\mathrm{cm}$ $60~{}\mathrm{cm}$ $4~{}\mathrm{cm}$ $1.5~{}\mathrm{cm}$ $0.02$
As detailed in Sec. III-B, Tac-Man exploits the compliance of the elastic deformable layer on the gripper to guide manipulation. This deformation provides critical tactile feedback by informing the gripper-object contact, which Tac-Man utilizes to adjust the gripper to maintain stable contact. Following previous studies [85], we employ rigid-body simulations to efficiently and rigorously evaluate our method on a wide range of objects. Despite their limitation in handling deformable materials, we do not use soft-body materials due to much lower speed and higher computational demands. To closely replicate the tactile feedback mechanism under these constraints in the rigid-body simulation, we devised specific adaptations to estimate the behavior of a deformable layer, as detailed below. ()$\prescript{\mathrm{sim}}{}{(}\cdot)$ denotes simulated variables.
Simulating sensor markers
We simulate the positions of the markers of the tactile sensors with point cloud grids. Specifically, we affix a point cloud grid, $\prescript{\mathrm{sim}}{}{\mathcal{C}}_{0}=\{\boldsymbol{p}^{0}_{i}|i=1,\dots,m\}$, on the grasping area of each gripper fingertip. We also add Gaussian noise $\gamma\sim\mathcal{N}(0,\zeta^{2})$ to the point cloud to imitate real-world sensor noise.
Simulating contact deviation
Initially, the gripper stably grasps the handle, and the initial contact set $\prescript{\mathrm{sim}}{}{\mathcal{C}_{1}}$ is determined as the markers close enough to the handle:
$\prescript{\mathrm{sim}}{}{\mathcal{C}_{1}}=\{\boldsymbol{p}^{1}|d(\boldsymbol{p}^{1})<\prescript{\mathrm{sim}}{}{\epsilon}\},$ (A1)
where $d(\boldsymbol{p})$ denotes the distance from point $\boldsymbol{p}$ to the handle. This contact set is then transformed to the handle’s frame $\{h\}$ and affixed to it, denoted as $\prescript{\mathrm{sim}}{}{\mathcal{C}_{h_{1}}}$. At each time step $i$ during manipulation, the contact set $\prescript{\mathrm{sim}}{}{\mathcal{C}_{i}}$ comprises the coordinated of the points affixed to the handle relative to the current gripper frame:
$\prescript{\mathrm{sim}}{}{\mathcal{C}_{i}}=\left\{\boldsymbol{T}_{h_{i}}^{i}\boldsymbol{p}^{h_{i}}+\begin{bmatrix}\gamma&\gamma&\gamma&0\end{bmatrix}^{\mathsf{T}}\mid\boldsymbol{p}^{h_{i}}\in\prescript{\mathrm{sim}}{}{\mathcal{C}_{h_{i}}}\right\}.$ (A2)
@@
The discrepancy between $\prescript{\mathrm{sim}}{}{\mathcal{C}_{1}}$ and $\prescript{\mathrm{sim}}{}{\mathcal{C}_{i}}$ in the gripper’s frame quantifies the contact deviation under the non-sliding stable contact conditions (Eq. 2). Based on this information, the iterative process between the execution and recovery stages in the simulation adheres to the methodology detailed in Sec. III and utilizes the same hyper-parameters as in the real-world experiments (Tab. I).
Emulating the hypothetic deformable layer
In rigid body simulations, we cannot simulate the elastic deformable layer, which causes misaligned dynamics compared to real-world scenarios. The fingers’ movement directly applies frictional forces to the object handle to move it, rather than through the contact between the layer and the handle. This can lead to unrealistic movements of the object part during the recovery stage, where Eq. 6 does not guarantee a successful recovery. To address this discrepancy, we have empirically increased the damping within the joint settings during this stage to emulate the effect of the deformable layer. This modification suppresses the unintended object movement, aligning the simulated interactions more closely with those observed in real-world conditions. Although these adaptations do not perfectly mirror real-world trajectories, they provide a reasonable framework to assess how our algorithm tracks and maintains contact stability, thereby validating our approach.
Determining hyper-parameters
Key hyper-parameters play a significant role in our simulation studies, and their values are determined based on empirical observations and testing, which we report in Tab. A1 to ensure transparency and reproducibility of the results.
-B Preprocess PartNet-Mobility
@@
The PartNet-Mobility dataset [36] comprises $1045$ articulated objects, and GAPartNet [38] enriches it with detailed part-level annotations. In this dataset, each object includes a base link and one or more moving links, connected via prismatic or revolute joints. Objects with multiple joints often necessitate sequential manipulations to fully access all their links (e.g. opening an oven door before pulling out the grill inside). For simplicity, we select one accessible handle on each of them for manipulation.
-C Generate Random 1-DoF Playboards
@@
Our simulation setup incorporates playboards consisting of a baseboard (white) and a toy train (red), and the train is restricted to a predetermined trajectory (grey). To achieve this, we constructed a groove on a cuboid board and set up a distance joint between the board and the train, limiting its movement to within the plane of the board. Two cylindrical locating pins beneath the train’s collision model are inserted into the groove to guide the train along the trajectory. The trajectories are sampled as Bézier curves, $B(s),s\in[0,1]$, with $n_{\mathrm{ctrl}}$ control points uniformly sampled and scaled to fit an $H\mathrm{cm}\times W\mathrm{cm}$ area, resulting in $(n_{\mathrm{ctrl}}-1)$-order curves. Any self-intersecting curves, taking into account a padding of $w$, are discarded. The groove is carved into a $(H+2w)~{}\mathrm{cm}\times(W+2w)~{}\mathrm{cm}$ board following the curve. The initial position of the train is set at $B(\eta)$ with a small positive $\eta$ to prevent initial pin-border collisions. The hyper-parameters for this setup are listed in Tab. A1.